OpenStudy (anonymous):

6. If 50% of a radioactive element remains after 5,000 years, what is its half-life?

OpenStudy (anonymous):

I really don't even understsand this at all.

OpenStudy (anonymous):

56

OpenStudy (anonymous):

@plm how did you get this?

OpenStudy (sleepyjess):

@plm Please do not give out direct answers.

OpenStudy (anonymous):

t1/2 = t * ln(2)/ln(N0/Nt) N0 = initial substance ; Nt = substance after (t) time t= time that passed

OpenStudy (anonymous):

t1/2 is the half time;

OpenStudy (anonymous):

the half time is 5000 years , Because you save after 5000 years 50% remains wich mean that 5000 years is the half time ? (I'm Not SURE) @goldennnn

OpenStudy (mathmate):

@borak @goldennnn Hint: "The radioactive half-life for a given radioisotope is the time for half the radioactive nuclei in any sample to undergo radioactive decay." from: http://hyperphysics.phy-astr.gsu.edu/hbase/nuclear/halfli.html

OpenStudy (anonymous):

i just seen this @borak @mathmate thank you both

OpenStudy (mathmate):

You're welcome! :)